# Utilizing mask-guided cross-image attention for zero-shot in-silico histopathologic image generation with a diffusion model

### v2

@@ Abstract
Utilizing generative models to create in-silico data presents an economical alternative to the traditional methods of staining, imaging, and annotating images in the computational pathology workflow. Specifically, appearance transfer diffusion models enable the generation of images without the need for model training, making the process swift and efficient. While originally developed for natural images, these models can transfer foreground objects from a source to a target domain, with less emphasis on the background. In computational pathology, however, every aspect of an image, including the background, can be crucial for comprehending the tumor micro-environment. In this study, we adapted an appearance transfer diffusion model to align with the demands of computational pathology by adjusting the AdaIN feature statistics in the denoising process. The effectiveness of this modified method was showcased through its application to epithelium segmentation. The results demonstrated superior performance compared to the baseline approach, indicating that the number of manual annotations necessary for model training could be reduced by 75% without sacrificing performance. The authors expect that this research will encourage the utilization of zero-shot diffusion models for the analysis of pathology images.
@@ 1 Introduction
Fig. 1: a Swapping key, values and queries from the appearance image (${i_{source}}$) with the target structure image (${m_{target}}$) we transfer the appearance of an image to the structure of a mask, generating an image (${i_{target}}$) matching a target mask (${m_{target}}$). b Generated images with matching masks.
@@
Deep learning model development in computational pathology requires datasets that contain meaningful, relevant, and accurate information. This enables models to learn how to perform tasks such as region segmentation or object detection. The creation of these datasets requires multiple steps such as sample preparation, staining, imaging, digitalization, quality control and manual annotations by a domain expert [1]. The manual annotation of large datasets with pixel-level accuracy is required for the training of supervised deep learning models but is a laborious and expensive task. In-silico data on the other hand is efficient and cheap to generate. Data with required properties can be generated through controllable parameterization [2]. Image generation has a tremendous impact on computational pathology providing methods to create datasets of high realism and fidelity [2]. This alleviates costly and time-consuming manual annotations [3, 4]. Applications of diffusion models in medicine, biology and pathology [2] span from segmentation [5], cell shape prediction [6], to artefact recovery [7]. Diffusion Probabilistic Models (DDPMs) are trained by iteratively learning how to remove added noise from an image. The denoising process is repeated at inference time to generate an in-silico image from gaussian noise. Diffusion models have been used to generate in-silico hematoxylin and eosin (H&E) stained images [7, 8] of rare cancer types that could not be differentiated from real images by pathologists, as well as for the synthesis of nuclei and classification datasets [9, 10, 11]. Text-conditioned diffusion models have been used to generate pathology in-silico H&E images from pathological reports [12] and for generating in-silico images at a gigapixel scale [13, 14]. We focus on the analysis of immunohistochemistry (IHC) stained images, which are characterized by an even larger diversity and complexity of visual patterns than H&E images. The recent work of Alaluf and al. shows that the appearance between images can be transferred during the desnoising diffusion process using cross- and self-attention mechanisms [15]. The keys, values and queries are first projected through the learned linear projections in each self-attention layer of the network. The attention scores, which indicate the relevance of each key to the corresponding queries, are then calculated for all keys and weighted by a Softmax operation. The aggregation of the scores results in an attention map which captures correspondences across the entire image. The queries capture the semantic information of each spatial location. The keys offer context for the queries. The values contain the content that is transferred from the source image to the target structure. In the typical case of appearance transfer of natural images, the foreground object (e.g. an animal) is of high importance in opposition to the background (e.g. the sky). However, in histopathology, both the foreground (e.g. epithelium) and the background (e.g. stroma) regions are important. Achieving a high level of realism for both is decisive to solve downstream tasks such as segmentation. We introduce in this work a novel appearance transfer diffusion models for generating in-silico data of IHC images in a zero-shot approach. By transferring the appearance of an existing image to an unpaired mask with realistic morphological structures, we generate realistic in-silico labeled images. Considering an image displaying regions of epithelium and stroma in each spatial organization, we seek to transfer the epithelium and stroma appearance into a different spatial organization while preserving the unique morphological characteristics of the two tissue types.
@@
Fig. 2: a Our model generates more realistic images than the baseline model. b From 2864 manually annotated images, we use a subset of 25% (N=716 images) and generate four in-silico images from each image. Semantic epithelium segmentation models trained on images generated with our diffusion model (red) perform better than trained on images generated with the baseline diffusion model (petrol). Upper (blue) and lower (black) baseline segmentation models are trained on all and the 25% subset of the labeled patches. Our test-set was annotated by 3 independent pathologists, mean dice score and confidence intervals (whiskers) are displayed. c Violin plot displaying the dice scores for each of the 43 FOVs and the three pathologists, showing significant differences (p $<$ 0.05, Wilcoxon signed-rank test, N=129) indicated with a star.
@@ 2 Methods
We adapted the baseline model developed by Alaluf et al. [15] to facilitate appearance transfer from existing IHC images to masks without the need for a classifier. This involved generating in-silico images with realistic IHC appearance and a corresponding pixel-precise segmentation mask based on existing true IHC images. The resulting augmented in-silico dataset was then utilized to train a downstream semantic segmentation model. The paper is structured into two main sections. The first section revisits the single class guided diffusion model as initially presented by Alaluf et al. [15]. The second section provides a detailed account of our technical contributions, particularly focusing on the proposed modifications of the original approach [15] to accommodate a multi-class setup.
@@ 2.1 Single-Class in-silico Image Generation
The guided diffusion model [15] requires a pair of input consisting of an image ${i_{source}}$ and a corresponding segmentation mask ${m_{source}}$ from the source domain as well as a segmentation mask ${m_{target}}$ from the target domain. This model then generates an in-silico image ${i_{target}}$ in the target domain that spatially corresponds to the target segmentation mask ${m_{target}}$ (Fig. 1). To achieve this, Stable Diffusion [16] a pretrained text-to-image model, is utilized for zero-shot appearance transfer (Fig. 1). By replacing the keys and values of the image ${i_{source}}$ while preserving the queries from the mask ${m_{target}}$, the model enables the transfer of the appearance contained in ${i_{source}}$ to the structure of ${m_{target}}$ between semantically similar objects. The initial step in the appearance transfer process involves calculating the inverted $z_{T}^{i_{source}}$ and $z_{T}^{m_{target}}$. Subsequently, at each forward step, these latents are passed to the denoising U-Net. Notably, the standard self-attention in the decoder of the U-Net is replaced with cross-image attention. During each denoising step t, two forward steps through the network are undertaken. The first step involves using the cross-attention layer: $\epsilon^{cross}=\epsilon_{\theta}^{cross}(z_{t}^{out})$. The second uses the self-attention layer of the network: $\epsilon^{self}=\epsilon^{self}(z_{t}^{out})$. The two noise predictions are finally aggregated: $\epsilon^{t}=\epsilon^{self}+\alpha(\epsilon^{cross}-\epsilon^{self})$. While ${\alpha}$ denotes the guidance scale, $z_{t-1}^{out}$ is the latent code at $t-1$ and ${\epsilon}^{t}$ is the modified noise. Additionally, the adaptive instance normalization (AdaIN) operation is employed to align the mean and variance of content features with those of the style features [17], thereby matching feature statistics between the latent representations.
@@ 2.2 Multi-Class in-silico Image Generation
We have made modifications to the appearance transfer mechanism. Initially, we introduce a class-specific feature statistics matching within the AdaIN operation, which is guided by the existing segmentation masks ${m_{c,source}}$ and ${m_{c,target}}$ through the class-specific masked latents, where c denotes the class. We applied a mask over the latents: $z_{t,c}^{i_{source}}=mask(z_{t}^{i_{source}})$ and $z_{t,c}^{m_{target}}=mask(z_{t}^{m_{target}})$ followed by the application of the AdaIN operation: $z_{t,c}^{out}=AdaIN(z_{t,c}^{m_{target}},z_{t,c}^{i_{source}})$ to align the color distributions of each class between the generated image and the input image. Additionally, we mapped the number of classes to the timesteps and alternately switched at each forward step of the denoising process between the class-specific masking of the AdaIN operation. For instance, in a two-class appearance transfer scenario, the mask of the foreground class was used in every even denoising step, while the mask of the background class was used in every odd denoising step. These two novel modifications of the appearance transfer mechanism significantly enhanced the fidelity and realism of the background region in the generated in-silico images compared to the original method [15].
Fig. 3: Ablation study to investigate performance saturation on 10% ($N=286images$) and 25% ($N=716images$) subsets. Dice scores of the segmented epithelium regions against three pathologist annotations (triangle, dotted lines) are displayed together with the corresponding mean (solid line) and standard deviation (colored area).
@@ 4 Results
In a series of quantitative experiments, we assessed the usefulness of the generated in-silico data by employing it in a semantic epithelium segmentation task. The mean dice scores, along with the associated low and high confidence intervals obtained by bootstrapping, the dice score of each FOV and each pathologist 999 times are reported as follows:
@@
Wilcoxon signed rank tests [19] revealed a significant difference (p=0.038) between the lower baseline model (b) and the model with augmented data (c), but not between the latter and the upper baseline model (a) (p=0.23). Additionally, Wilcoxon signed rank tests show a significant difference (p=0.04) between the lower baseline model (b) and the baseline model [6] including augmented data (d) and between the latter (d) and the upper baseline model (a) (p=$2.8\times 10^{-5}$). Furthermore, we find that each fold of added data from our model yields significant improvements over the lower baseline segmentation model (p$<$0.05, Wilcoxon signed-rank tests, 2c). The use of the proposed diffusion model demonstrated an improvement over the baseline diffusion model for data augmentation. Performance achieved with the latter saturated after one-fold of in-silico data generation, while further performance gains were observed with the proposed model (see Fig. 2b,c). The results of an ablation study, where the dataset was reduced to 25% and 10% of its original size (Fig. 3), illustrated the impact of the number of augmentation folds on performance. The segmentation accuracy saturated after four and eight augmentation folds, respectively. In a series of qualitative experiments, we assessed the accuracy of the generated IHC images with a board-certified pathologist. Our objective was to identify potential areas for improvement. While the pathologist found that most of the morphologies were realistically generated, some differences were observed (Fig. 4). These variances included overcrowded nuclear regions displaying unrealistic morphologies for NSCLC, resembling a papillary pattern without the supportive stroma. Additionally, discrepancies were noted in regions with exceedingly high nuclear density, lacking corresponding cell membranes, and exhibiting atypical representation of nuclear size and shape variation (anisokaryosis).
@@ 4.1 Discussion
In this study, we present a novel diffusion model-based data generation pipeline tailored to meet the needs of computational pathology. This adapted model, originally trained on natural images, is applied in a zero-shot setting. By extending a baseline diffusion model to a multi-class setting, we can produce realistic IHC images and reduce the dependency on labeled images by approximately 75% without compromising the performance of downstream segmentation tasks. Collaborating with a certified pathologist, we identified that real and in-silico images are primarily distinguishable through nuclear morphology, prompting the consideration of training a histopathology-specific backbone to replace the current model. We acknowledge that our approach requires segmentation masks and is not applicable in a fully unsupervised setup that is devoid of annotations. There is a need to further explore the performance of our model in scenarios involving more than two classes in an image or with inaccurate annotations. We believe this study demonstrates the efficient utilization of zero-shot models for meaningful in-silico data generation in computational pathology.

### v3

@@ Abstract
Utilizing generative models to create in-silico data presents an economical alternative to the traditional methods of staining, imaging, and annotating images in the computational pathology workflow. Specifically, appearance transfer diffusion models enable the generation of images without the need for model training, making the process swift and efficient. While originally developed for natural images, these models can transfer foreground objects from a source to a target domain, with less emphasis on the background. In computational pathology, however, every aspect of an image, including the background, can be crucial for understanding the tumor micro-environment. In this study, we adapted an appearance transfer diffusion model to align with the demands of computational pathology by adjusting the AdaIN feature statistics in the denoising process. The effectiveness of this modified method was demonstrated through its application to epithelium segmentation. The results demonstrated superior performance compared to the baseline approach, indicating that the number of manual annotations necessary for model training could be reduced by 75% without sacrificing accuracy. The authors expect that this research will promote the use of zero-shot diffusion models in computational pathology.
@@ 1 Introduction
Fig. 1: a By swapping key, values and queries from the appearance image (${i_{source}}$) with the target structure image (${m_{target}}$) we transfer the appearance of an image to the structure of a mask, generating an image (${i_{target}}$) that matches the target mask (${m_{target}}$). This process alternates AdaIN feature statistics matching across different classes, utilizing the existing segmentation masks on the latent representations ${z^{source}}$ and ${z^{target}}$. b Real images and their corresponding in-silico images, which share the same appearance but different structures. c A UMAP plot reveals overlapping clusters of real and in-silico images.
@@
Developing deep learning models in computational pathology requires datasets with accurate, relevant information for tasks like segmentation and object detection. Creating these datasets involves stages such as sample preparation, staining, imaging, and expert manual annotation, which is labor-intensive and costly [1]. In contrast, in-silico data provides a more efficient, cost-effective alternative by generating datasets with the necessary properties through controllable parameterization [2]. Image generation significantly impacts computational pathology by enabling the creation of realistic datasets, reducing the need for costly manual annotations [2, 3, 4]. Diffusion models in medicine, biology, and pathology [2] have diverse applications, including segmentation [5], cell shape prediction [6], and artifact recovery [7]. These models, trained to remove noise from images, can generate in-silico images from Gaussian noise. They have been used to create realistic H&E stained images [7, 8] and datasets for rare cancers and nuclei classification [9, 10, 11, 12]. Text-conditioned diffusion models can generate pathology images from reports [13] or at gigapixel scales [14, 15]. Our work focuses on analyzing the greater complexity of immunohistochemistry (IHC) stained images. The recent work of Alaluf and al. shows that the appearance between images can be transferred during the desnoising diffusion process using cross- and self-attention mechanisms [16]. The keys, values and queries are first projected through the learned linear projections in each self-attention layer of the network. The attention scores, which indicate the relevance of each key to the corresponding queries, are then calculated for all keys and weighted by a Softmax operation. The aggregation of the scores results in an attention map which captures correspondences across the entire image. The queries capture the semantic information of each spatial location. The keys offer context for the queries. The values contain the content that is transferred from the source image to the target structure. In the case of appearance transfer for natural images, the foreground object (e.g. an animal) is of high importance in opposition to the background (e.g. the sky). However, in histopathology, both the foreground (e.g. epithelium) and the background (e.g. stroma) regions are important. Achieving a high level of realism for both is decisive to solve downstream tasks such as segmentation. In this study, we introduce a novel appearance transfer diffusion models for generating in-silico data of IHC images in a zero-shot approach. By transferring the appearance of an existing image to an unpaired mask with realistic morphological structures, we generate images with high fidelity. Specifically, we aim to transfer the appearance of epithelium and stroma into different spatial organizations, while preserving the unique morphological characteristics of each tissue type. This method has the potential to generate highly realistic and diverse datasets for training and evaluating computational models in pathology.
@@
Fig. 2: a Our model generates more realistic images than the baseline model. b From a dataset of 2,864 manually annotated images, we use a 25% subset (N=716 images) and generate four in-silico images from each image. Semantic epithelium segmentation models trained on images generated with our diffusion model (red) outperform those trained on images generated by the baseline model (petrol). The upper (blue) and lower (black) baseline segmentation models were trained on the full dataset and the 25% subset of the labeled patches, respectively. Our test-set was annotated by three independent pathologists. The mean Dice score and corresponding confidence intervals (whiskers) are displayed. c A violin plot showing Dice scores for each of the 43 FOVs and the three pathologists reveals significant differences (p $<$ 0.05, Wilcoxon signed-rank test, N=129), with statistical significance indicated by a star.
@@ 2 Methods
We adapted the baseline model developed by Alaluf et al. [16] to enable appearance transfer from existing IHC images to masks. This adaptation involved generating in-silico images with realistic IHC appearance and corresponding pixel-accurate segmentation masks, derived from real IHC images. The augmented in-silico dataset was then used to train a downstream semantic segmentation model. The paper is organized into two main sections. he first section revisits the single-class guided diffusion model originally presented by Alaluf et al. [16]. The second section provides a detailed description of our technical contributions, with a particular focus on the modifications made to the original approach [16] to support a multi-class setup.
@@ 2.1 Single-Class in-silico Image Generation
The guided diffusion model [16] requires a pair of inputs: an image ${i_{source}}$ and its corresponding segmentation mask ${m_{source}}$ from the source domain, along with a segmentation mask ${m_{target}}$ from the target domain. The model then generates an in-silico image ${i_{target}}$ in the target domain that spatially corresponds to the target segmentation mask ${m_{target}}$ (Fig. 1). To achieve this, Stable Diffusion [17] a pretrained text-to-image model, is utilized for zero-shot appearance transfer (Fig. 1). By replacing the keys and values of the image ${i_{source}}$ while preserving the queries from the mask ${m_{target}}$, the model enables the transfer of the appearance contained in ${i_{source}}$ to the structure of ${m_{target}}$ between semantically similar objects. The appearance transfer process begins with the calculation of the inverted latents $z_{T}^{i_{source}}$ and $z_{T}^{m_{target}}$. These latents are then passed through the denoising U-Net at each forward step. Notably, the standard self-attention in the U-Net decoder is replaced with cross-image attention. During each denoising step t, two forward passes are performed. The first uses the cross-attention layer: $\epsilon^{cross}=\epsilon_{\theta}^{cross}(z_{t}^{out})$. The second step uses the self-attention layer: $\epsilon^{self}=\epsilon^{self}(z_{t}^{out})$. The two noise predictions are then combined as follows: $\epsilon^{t}=\epsilon^{self}+\alpha(\epsilon^{cross}-\epsilon^{self})$. Where ${\alpha}$ represents the guidance scale, $z_{t-1}^{out}$ is the latent code at step $t-1$ and ${\epsilon}^{t}$ is the modified noise. Additionally, the Adaptive Instance Normalization (AdaIN) operation is used to align the mean and variance of content features with those of style features [18], thereby matching the feature statistics between the latent representations.
@@ 2.2 Multi-Class in-silico Image Generation
We have introduced modifications to the appearance transfer mechanism. First, we incorporated class-specific feature statistics matching within the AdaIN operation, guided by the existing segmentation masks ${m_{c,source}}$ and ${m_{c,target}}$ through class-specific masked latents, where c denotes the class. To achieve this, we applied a mask to the latents: $z_{t,c}^{i_{source}}=mask(z_{t}^{i_{source}})$ and $z_{t,c}^{m_{target}}=mask(z_{t}^{m_{target}})$ followed by the AdaIN operation: $z_{t,c}^{out}=AdaIN(z_{t,c}^{m_{target}},z_{t,c}^{i_{source}})$ which aligns the color distributions for each class between the generated and input images. Additionally, we mapped the class labels to the timesteps, alternating between class-specific AdaIN masking at each forward step of the denoising process. For instance, in a two-class appearance transfer scenario, the foreground class mask was applied during every even denoising step, while the background class mask was applied during every odd denoising step. These two novel modifications significantly enhanced the fidelity and realism of the background region in the generated in-silico images, as compared to the original method [16].
Fig. 3: An ablation study was conducted to investigate performance saturation using 10% ($N=286$ images) and 25% ($N=716$ images) subsets. The Dice scores for the segmented epithelium regions, compared against annotations from three pathologists (denoted by triangles and dotted lines), are displayed alongside the corresponding mean (solid line) and standard deviation (shaded area).
@@ 4 Results
In a series of quantitative experiments, we evaluated the utility of the generated in-silico data by applying it to a semantic epithelium segmentation task. The mean Dice scores, along with the associated low and high confidence intervals obtained through bootstrapping (999 resamples), are reported as follows:
@@
Wilcoxon signed rank tests [20] revealed a significant difference (p=0.038) between the lower baseline model (b) and the model with augmented data (c), but not between the latter and the upper baseline model (a) (p=0.23). Additionally, Wilcoxon signed rank tests show a significant difference (p=0.04) between the lower baseline model (b) and the baseline model [6] including augmented data (d) and between the latter (d) and the upper baseline model (a) (p=$2.8\times 10^{-5}$). Furthermore, we find that each fold of added data from our model yields significant improvements over the lower baseline segmentation model (p$<$0.05, Wilcoxon signed-rank tests, 2c). The proposed diffusion model demonstrated an improvement over the baseline diffusion model. Performance achieved with the latter saturated after one-fold of in-silico data generation, while further performance gains were observed with the proposed model (see Fig. 2b,c). An ablation study, in which the dataset was reduced to 25% and 10% of its original size (Fig. 3), demonstrated the impact of the number of augmentation folds on segmentation performance. Accuracy saturated after four and eight folds of augmentation, respectively. In a series of qualitative experiments, we assessed the accuracy of the generated IHC images with a board-certified pathologist. The goal was to identify potential areas for improvement. While the pathologist found that most morphologies were realistically generated, some discrepancies were noted (Fig. 4). These included overcrowded nuclear regions with unrealistic morphologies for NSCLC, resembling a papillary pattern without the corresponding stroma. Additionally, regions with extremely high nuclear density lacked cell membranes and exhibited atypical variations in nuclear size and shape (anisokaryosis).
@@ 4.1 Discussion
In this study, we introduce a novel diffusion model-based data generation pipeline optimized for computational pathology. The model, originally trained on natural images, is adapted and applied in a zero-shot setting. By extending a baseline diffusion model to a multi-class framework, we can generate realistic IHC images, reducing reliance on labeled data by approximately 75% without compromising performance in downstream segmentation tasks. In collaboration with a certified pathologist, we found that real and in-silico images are primarily distinguishable based on nuclear morphology. This observation suggests the potential for training a histopathology-specific backbone to improve model performance. However, we acknowledge that our approach requires segmentation masks and is not fully applicable in a fully unsupervised setup without annotations. Further investigation is required to assess the model’s performance in scenarios involving more than two classes or inaccurate annotations. Additionally, we plan to compare our approach with the Semantic Diffusion Model [21] to evaluate its relative strengths. We also aim to explore how in-silico data can help mitigate performance gaps in deep learning models for computational pathology, especially in cases of imbalanced datasets [22]. We believe this study highlights the effective use of zero-shot models for generating high-quality in-silico data in computational pathology.